OpenStudy (anonymous):

Find the slope of the line passing through the points (0, 4) and (–8, –1)

OpenStudy (amistre64):

we can move the set so that one of them is at the origin ... i would suggest the 0,4 (0, 4) and (–8, –1) +4 +4 ------------------ 0,0 -8, 3 now we can easily read the slope from this point as: y/x

OpenStudy (amistre64):

lol, forgot how to subtract ...

OpenStudy (amistre64):

0, 4) and (–8, –1) -4 -4 ------------------ 0,0 -8, -5

OpenStudy (anonymous):

slope formula: m = [y1 - y2] / [x1 - x2] m = (4+1)/ (0+8) = 5/8

OpenStudy (anonymous):

slope is 5/8